Paris Manufacturing that produces computer hard drives has had problems with hard drive defects in the past. The relevant information on their current inspection system follows.

Current Manual Inspection System
Annual fixed cost = $42,000
Inspection variable cost per unit = $12.50

A marketing representative of NEW-SPEC, a firm that specializes in providing manual inspection processes for other firms, approached Paris Manufacturing and offered to inspect parts for $20.50 each with no fixed cost. It assured Paris Manufacturing that the accuracy and quality of its manual inspections would equal that of the automated inspection system. Demand for the upcoming year is forecast to be 11,000 units. Should the manufacturer accept the offer? This problem includes break-even calculations. For a review of how to calculate break-even analysis, please refer to Chapter 6 – Equations 6.1, 6.2, and 6.3. Round your answer for the breakeven volume to the nearest whole number and round your answer for the amount of saving/loss to the nearest dollar.

Breakeven volume:______ hard drives

Since the annual demand of 11,000 hard drives is -Select-greater/less than the breakeven volume, the firm -Select-should/should not accept the offer. They will -Select-save/loseI $_______by outsourcing.
Answer #1

buy price= 20.5 per inspection

Make at Paris manufacturing

Fixed cost= 42,000

Variable cost= 12.5 per inspec

Profit of making inhouse= 20.5-12.5= 8 per inspec

The total unit produced is x, then the total profit will be Y

Total profit= Profit per piece*x - Fixed cost

Y= 8*x – 42000

For break even, total profit y= 0

8*x= 42000

X= 6,000 units (Answer a)

If demand for hard drives is, x= 11,000 units

Answer b: they should not accept the offer, manufacture in the house as the demand is more than breakeven quantity

Total outsourcing cost of 11,000 inits= 11000*20.5 = $225,500

Total inhouse cost of 11,000 inits= 11000*12.5 + 42000= $54,500

difference in cost= $171000

they may lose $171,000 by outsourcing

Please ask, if you have any doubts through the comment section. Do rate the answer